OpenStudy (anonymous):

If y = –49 when x = –7, find x when y = –7. a. y = 7x; –1 c. y = –7x; 1 b. y = 7x; –0.9 d. y = 6x; –1

OpenStudy (ajprincess):

Let y=kx. When y=-49 and x=-7 -49=k(-7) find k. when y=-7 y=kx -7=kx. plug in the value of k and find the value of x. Does that help? @yololife1423

OpenStudy (anonymous):

yhja thanks @ajprincess

OpenStudy (ajprincess):

welcome:)